OpenStudy (anonymous):

State the domain of the numbers displayed in column x and column y. x y 1 5 2 7 3 9

OpenStudy (anonymous):

test conistency and if possible solve x+2y+3z=14 3x+y+2z=11 2x+3y+z=11

OpenStudy (anonymous):

all the xs

OpenStudy (anonymous):

give me the solution